OpenStudy (bayleejadee):

A farmer grows two types of grain in his fields. He has a seed budget of $17,000.00 and has spent $9,200.00 to plant wheat seed so far. If barley seed costs $41.00 per acre to plant, what is the maximum number of acres of barley the farmer can plant and still be under budget? 190 acres 414 acres 191 acres 224 acres

zepdrix (zepdrix):

Let the number of unknown acres be x. Then if each acre costs 41 to seed, then x acres will cost 41x. This is the cost to seed x acres of farm: 41x. Let's add to this, the amount he's already spent: 9200. 41x+9200 This amount must not exceed farmer Billy's budget: 17000. So we're saying this amount must be `equal or less than` the budget amount. \(\large\rm 41x+9200\le17000\)

zepdrix (zepdrix):

Solve for x. It has 9200 being added, so we'll undo that by subtracting 9200 from each side,\[\large\rm 41x+9200-9200\le17000-9200\]\[\large\rm 41x\le7800\]

zepdrix (zepdrix):

What next? 41 is `multiplying` the x. How do we undo multiplication? Hmm any ideas?

OpenStudy (bayleejadee):

@zepdrix with division?

zepdrix (zepdrix):

Yes, division is the opposite of multiplication. That seems like a good idea. So we divide by 41,\[\large\rm \frac{\cancel{41}x}{\cancel{41}}\le\frac{7800}{41}\]

zepdrix (zepdrix):

\[\large\rm x\le 190.244\]

zepdrix (zepdrix):

So farmer Billy Bob can afford to seed 190.244 acres, but no more than that. So which option should he choose?

OpenStudy (bayleejadee):

190 acres (:

zepdrix (zepdrix):

because 191 would be too much? Yayy good job \c:/